### v4

# SKEL-CF: Coarse-to-Fine Biomechanical Skeleton and Surface Mesh Recovery

Abstract
Parametric 3D human models such as SMPL have driven significant advances in human pose and shape estimation, yet their simplified kinematics limit biomechanical realism. The recently proposed SKEL model addresses this limitation by re-rigging SMPL with an anatomically accurate skeleton. However, estimating SKEL parameters directly remains challenging due to limited training data, perspective ambiguities, and the inherent complexity of human articulation. We introduce SKEL-CF, a coarse-to-fine framework for SKEL parameter estimation. SKEL-CF employs a transformer-based encoder–decoder architecture, where the encoder predicts coarse camera and SKEL parameters, and the decoder progressively refines them in successive layers. To ensure anatomically consistent supervision, we convert the existing SMPL-based dataset 4DHuman into a SKEL-aligned version, 4DHuman-SKEL, providing high-quality training data for SKEL estimation. In addition, to mitigate depth and scale ambiguities, we explicitly incorporate camera modeling into the SKEL-CF pipeline and demonstrate its importance across diverse viewpoints. Extensive experiments validate the effectiveness of the proposed design. On the challenging MOYO dataset, SKEL-CF achieves 85.0 MPJPE / 51.4 PA-MPJPE, significantly outperforming the previous SKEL-based state-of-the-art HSMR (104.5 / 79.6). These results establish SKEL-CF as a scalable and anatomically faithful framework for human motion analysis, facilitating the use of computer vision techniques in biomechanics-related analysis. Our implementation is available on the project page: https://pokerman8.github.io/SKEL-CF/.
Figure 1: Visualization of SKEL-CF on web images. Compared with SMPL-based state-of-the-art CameraHMR [27], our SKEL-based model produces more natural joint motions (see the side-view zoomed knee in blue). Compared with HSMR [39], the state-of-the-art SKEL-based method, SKEL-CF achieves more accurate skeleton and mesh reconstruction (see zoomed hand in orange).
1 Introduction
In recent years, 3D human pose and shape estimation has made remarkable progress, enabling a wide range of downstream applications [5, 8, 19, 37, 32]. However, adoption in biomechanics, a domain where such techniques could be particularly impactful, remains limited. This gap arises because existing parametric models, such as SMPL [23], SMPL-X [28], and GHUM [40], fall short of the stringent biomechanical requirements. Their simplified kinematics and loosely constrained axis–angle representations often compromise biomechanical realism, especially in complex articulations such as deep squats or yoga poses [39]. In these scenarios, the estimated parameters must correspond to anatomically accurate skeletons, respect joint limits, and ensure physically plausible motion with high precision. As illustrated in Figure 1 (a), even the state-of-the-art SMPL-based method CameraHMR [27] can produce anatomically implausible results, such as unnatural knee bending.
To address these limitations, the recently proposed SKEL [15] model redefines the foundation of human representation by re-rigging SMPL with an anatomically accurate skeleton and realistic joint constraints. This makes it a promising step toward bridging computer vision and biomechanics, offering a representation that is both visually coherent and biomechanically faithful. HSMR [39] is the first transformer-based method to reconstruct a human body in the SKEL space from a single image. It converts the predicted SMPL ground truth from existing datasets into SKEL pseudo ground truth for supervised training. However, estimating SKEL from a single image remains challenging, as it requires inferring 3D structure from inherently limited 2D observations. Consequently, the predictions often lack accuracy (see Figure 1 (b)), and empirical metrics such as MPJPE remain inferior to SMPL-based models.
In this work, we take a step forward in advancing SKEL-based human mesh recovery. We observe that estimating SKEL parameters in a coarse-to-fine manner significantly improves performance. Specifically, our transformer encoder first predicts coarse camera and SKEL parameters, while the decoder refines these estimates progressively, layer by layer, leading to more accurate and stable results. To address challenges arising from diverse camera viewpoints, we incorporate camera intrinsic estimation following CameraHMR [27], which enhances the robustness of the reconstructed mesh across varying perspectives. Moreover, by leveraging the high-quality SMPL estimations provided by the recent 4DHuman dataset, we generate large-scale, high-fidelity SKEL annotations to form the 4DHuman-SKEL dataset, enabling more reliable training for SKEL estimation.
Empirically, our proposed SKEL-CF achieves substantial performance gains across diverse benchmarks. On the challenging MOYO dataset [33], SKEL-CF reaches 85.0 MPJPE and 51.4 PA-MPJPE, significantly outperforming the previous SKEL-based state-of-the-art HSMR [39], which reports 104.5 and 79.6, respectively. Furthermore, SKEL-CF attains comparable or superior quantitative performance to leading SMPL-based approaches such as CameraHMR [27] across multiple datasets, including 3DPW [35], EMDB [14], and MOYO [33], while producing biomechanically faithful and visually consistent mesh reconstructions (See igure 1 (c)) These results position SKEL-CF as a robust and anatomically realistic solution for practical human mesh recovery in real-world scenarios. We also conduct ablation studies to validate the contribution of each proposed component in our framework.
To conclude, our key contributions are summarized as follows:
  • We propose SKEL-CF, a coarse-to-fine transformer framework that significantly advances SKEL-based human mesh recovery. By explicitly modeling the camera and refining SKEL parameters layer by layer, SKEL-CF achieves state-of-the-art performance across multiple benchmarks while maintaining biomechanical fidelity.
  • We demonstrate that estimating SKEL parameters in a coarse-to-fine and iterative refinement manner leads to substantial performance gains. The decoder progressively improves predictions across layers, effectively handling complex articulations and challenging poses.
  • We introduce 4DHuman-SKEL, a large-scale, high-quality dataset with anatomically consistent SKEL annotations converted from 4DHuman [6]. This dataset provides reliable supervision and unified skeleton–mesh alignment, enabling robust SKEL-based model training.
Our implementation is available on the project page: https://pokerman8.github.io/SKEL-CF/.
Figure 2: Overview of the proposed SKEL-CF. Our method estimates SKEL parameters from a single image using an encoder–decoder architecture that performs coarse-to-fine estimation. The encoder produces initial predictions of the camera extrinsics $\boldsymbol{\pi}$, shape parameters $\boldsymbol{\beta}$, and pose parameters $\boldsymbol{\theta}$. The decoder then refines these predictions progressively across layers. In addition, we adopt the pretrained camera model from CameraHMR [27] to estimate the camera intrinsics, and keep its parameters frozen throughout training.
2 Related Work
Parametric human body model.
Parametric human body models are widely used for 3D human shape and pose estimation, as they offer a structured low-dimensional representation of human geometry. SMPL [23] represents the human mesh with fixed topology, using low-dimensional shape and pose parameters with linear blend skinning and pose-dependent corrective shapes for realistic articulation. SMPL-H [29] augments this framework with articulated hands, while SMPL-X [28] further integrates the FLAME [20] facial model, yielding a unified representation of body, hands, and expressive face. While the SMPL family provides a compact and effective representation of the human body surface, it employs a simplified kinematic structure that deviates from the anatomical skeletal system of real humans, allowing anatomically implausible poses (e.g., the knee can exhibit unrealistically free rotation, see Figure 1 (b)). SKEL [15] introduces anatomically accurate joint definitions and a bone hierarchy embedded within a parametric human body model, overcoming limitations of simplified kinematic structures in models such as SMPL. It can produce a skeleton mesh aligned with the body surface that is suitable for applications requiring anatomically valid joint locations and realistic bone articulations, such as biomechanical analysis or as a basis for further physics-based simulation.
Human mesh recovery.
The lack of 3D annotations for in-the-wild images remains a major obstacle in Human mesh recovery. SMPLify [2] mitigates this by iteratively optimizing SMPL [23] parameters from 2D keypoints, yielding accurate results but at high computational cost. HMR [11] introduces the first end-to-end framework to predict 3D pose and shape from in-the-wild images using only 2D supervision, where adversarial learning with mocap data provides strong pose and shape priors to compensate for the missing 3D information. 4DHuman [6] constructs a large-scale dataset by fitting approximately 3 million images with ProHMR [18], generating pseudo-ground-truth (p-GT) 3D annotations alongside 2D keypoints, and further upgrades the network architecture from CNN to Vision Transformer (ViT). However, TokenHMR [4] identifies a mismatch between 3D and 2D keypoints, where enforcing 2D keypoint alignment can degrade 3D evaluation metrics, partly due to the use of fixed camera intrinsics and p-GT annotations in the 4DHuman [6] dataset. To address this, TokenHMR [4] introduces the TALS loss, which down-weights the 2D keypoint loss when the L2 distance between predicted and ground-truth keypoints falls within a predefined threshold, thus preventing overfitting to noisy p-GT annotations. In addition, it leverages a quantized token dictionary, constructed by pre-training a Vector Quantized-VAE (VQ-VAE) [34] on motion capture datasets [25, 33], to mitigate 3D pose ambiguity. CameraHMR [27] improves the 4DHuman [6] dataset through CamSMPLify, which refines body representations using dense keypoints to overcome the average body limitation. To tackle the fixed-camera constraint, it introduces HumanFOV [27], transforming weak-perspective projection into a fully perspective projection.
SMPL-to-SKEL dataset conversion.
The first large-scale 3D pseudo-ground-truth (p-GT) dataset was introduced by 4DHuman [6], which extends to unlabeled datasets such as InstaVariety [13], AVA [7], and AI Challenger [31] by generating p-GT annotations. For each image, an off-the-shelf detector [3] and a body keypoint estimator are applied to obtain bounding boxes and 2D keypoints. A SMPL [23] mesh is then fitted to these keypoints using ProHMR [18], producing pseudo-ground-truth SMPL parameters. Despite its scale and utility, the 4DHuman dataset still contains inaccuracies and artifacts. To that end, CameraHMR [27] refines the dataset using its proposed CamSMPLify method, alongside additional techniques to enhance annotation quality. While large-scale datasets of SMPL parameters have been constructed, there remains a need to develop datasets for SKEL [15] parameters. HSMR [39] addresses this by first converting the 4DHuman SMPL parameters to SKEL parameters using SKEL fitting [15]. To mitigate the inaccuracies in 4DHuman annotations, we further apply SKEL fitting to the refined SMPL annotations provided by CameraHMR.
3 Method
In this section, we introduce our approach, SKEL-CF, whose overall pipeline is illustrated in Figure 2. SKEL-CF aims to estimate SKEL parameters from a single image using an encoder–decoder architecture that performs coarse-to-fine estimation. The encoder produces initial predictions of the camera extrinsics $\boldsymbol{\pi_{0}}$, shape parameters $\boldsymbol{\beta_{0}}$, and pose parameters $\boldsymbol{\theta_{0}}$, while the decoder progressively refines these predictions across layers. To enhance robustness to camera variations, we adopt the pretrained camera model from CameraHMR [27] to estimate the camera intrinsics, and keep its parameters frozen throughout training.
This section is organized as follows. We first introduce the SKEL model and the 4DHuman-SKEL dataset that we construct to supervise SKEL learning in Section 3.1. We then describe the details of SKEL-CF and its implementation in Section 3.2.
3.1 Preliminaries
SKEL model.
The SKEL model [15] is a parametric representation of the human body that enables the joint reconstruction of both the skin surface and the skeletal mesh. Given the pose parameters $q\in\mathbb{R}^{46}$ and the shape parameters $\boldsymbol{\beta}\in\mathbb{R}^{10}$ (corresponding to the top 10 shape components), SKEL generates the human mesh $M\in\mathbb{R}^{6890\times 3}$ and the skeleton mesh $S_{k}\in\mathbb{R}^{24752\times 3}$. Unlike SMPL, which defines pose parameters $\theta\in\mathbb{R}^{72}$ using unrestricted three-degree-of-freedom ball joints, SKEL constrains the motion of each joint, reducing the parameter space from 72 to 46. This formulation ensures anatomically consistent and biomechanically valid poses [15, 39]. Consequently, SKEL produces physically plausible human motion and is particularly suitable for downstream applications such as motion analysis, rehabilitation, and human–robot interaction.
4DHuman-SKEL: a high-quality dataset for SKEL training.
HSMR [39] initially generated pseudo ground-truth (p-GT) SKEL parameters by converting the SMPL parameters into the SKEL parameter space. While this conversion provided a practical starting point, its quality was limited by the low-resolution images and noisy annotations in the original 4DHuman [6] dataset, which aggregates samples from Human3.6M [9], MPI-INF-3DHP [26], COCO [22], MPII [1], AI Challenger [38], AVA [7] and InstaVariety [12]. With the recent release of the refined 4DHuman SMPL annotations [6] from CameraHMR [27], we revisit this process to obtain higher-quality SKEL p-GT annotations. Consequently, we exclude AVA, as CameraHMR does not provide refined SMPL annotations for this subset, and denote the resulting dataset as 4DHuman-SKEL.
Specifically, we reconstruct the human mesh from the provided SMPL [23] parameters using the neutral SMPL model. Our overall optimization pipeline largely follows the SKEL [15] fitting protocol (i.e., starting from an initial set of SKEL parameters and performing iterative optimization with a hierarchical refinement schedule), but we introduce a small yet important adaptation to better align with the refined 4DHuman dataset [6]. Different from SKEL, since both models share identical global orientation and translation in our setting, we directly convert the SMPL global orientation from axis–angle to Euler angles to comply with the SKEL convention and fix these parameters during optimization. We observe that further optimizing or re-learning global orientation and translation does not improve alignment quality on the refined dataset. This process takes approximately 58 hours to process all 3M images on a single RTX 3090 GPU.
3.2 SKEL-CF
Methods 3DPW [35] Human3.6M [9] MOYO [33] MOYO-HARD
MPJPE $\downarrow$ PA-MPJPE $\downarrow$ MPJPE $\downarrow$ PA-MPJPE $\downarrow$ MPJPE $\downarrow$ PA-MPJPE $\downarrow$ MPJPE $\downarrow$ PA-MPJPE $\downarrow$
Two-stage Approaches
HMR2.0 + SKEL fit [39, 6] 81.0 54.4 53.6 34.1 130.5 93.7 - -
CameraHMR + SKEL fit [39, 27] 70.4 41.8 - - 75.5 49.9 88.7 61.4
End to End Approaches
HSMR [39] 81.5 54.8 50.4 32.9 104.5 79.6 120.0 97.7
SKEL-CF (Ours) 61.5 38.7 39.0 31.2 85.0 51.4 90.0 61.5
Table 1: Quantitative comparison of skeleton and surface mesh recovery using the SKEL [15] human model on 3DPW [35], Human3.6M [9], and MOYO [33]. The proposed SKEL-CF delivers consistent accuracy gains over previous end-to-end approach. denotes results reported by [39], denotes fitting the outputs of [27] into the SKEL [15] model.
Overall pipeline.
As illustrated in Figure 2, SKEL-CF employs an encoder–decoder architecture with a coarse-to-fine strategy to estimate SKEL parameters from a single RGB image. We denote the target parameter set as $\boldsymbol{\Theta}=\{\boldsymbol{\theta},\boldsymbol{\beta},\boldsymbol{\pi}\}$, where $\boldsymbol{\theta}$ represents the human pose, $\boldsymbol{\beta}$ the body shape, and $\boldsymbol{\pi}$ the camera extrinsics.
Given an input image, the encoder first encodes it into visual tokens and, together with camera intrinsic information, extracts contextual features through standard self-attention.
Using these features, the encoder produces initial estimates $\boldsymbol{\Theta}_{0}=\{\boldsymbol{\theta}_{0},\boldsymbol{\beta}_{0},\boldsymbol{\pi}_{0}\}$. These initial predictions are then progressively refined through $T$ decoder layers using cross-attention with the encoded visual features as context. The final output of SKEL-CF is $\boldsymbol{\Theta}_{T}=\{\boldsymbol{\theta}_{T},\boldsymbol{\beta}_{T},\boldsymbol{\pi}_{T}\}$.
SKEL learning objectives.
We learn the target parameters $\boldsymbol{\hat{\Theta}}=\{\boldsymbol{\hat{\theta}},\boldsymbol{\hat{\beta}},\boldsymbol{\hat{\pi}}\}$ using a loss integrating keypoint-level and parameter-level supervision:
$\mathcal{L}_{\text{skel}}=\underbrace{\lambda_{kp}\mathcal{L}_{\text{kp}}}_{\text{keypoint-level}}+\underbrace{\lambda_{\beta}\mathcal{L}_{\beta}+\lambda_{\theta}\mathcal{L}_{\theta}}_{\text{parameter-level}}.$ (1)
where $\lambda_{kp}$, $\lambda_{\beta}$, and $\lambda_{\theta}$ are hyperparameters controlling the relative weight of each term. This formulation allows us to simultaneously supervise the reconstructed 3D/2D joints and the underlying pose and shape parameters. Note that no direct camera extrinsic supervision $\boldsymbol{\hat{\pi}}$ is provided; the ability to predict camera extrinsics is only implicitly learned through weak 2D keypoint reprojection, making the training more challenging.
Keypoint-level supervision.
Given the pose $\boldsymbol{\theta}$ and shape $\boldsymbol{\beta}$, we first reconstruct 3D joint positions $\mathbf{J}_{3d}\in\mathbb{R}^{K\times 3}$ via forward kinematics, where $K$ is the number of joints. These 3D joints are projected onto the image plane using the predicted camera parameters to obtain 2D joint locations $\mathbf{J}_{2d}\in\mathbb{R}^{K\times 2}$. The coarse keypoint loss $\mathcal{L}_{\text{kp}}$ enforces consistency with pseudo-ground-truth 3D and 2D joints:
$\mathcal{L}_{\text{kp}}=\|\hat{\mathbf{J}}_{3d}-\mathbf{J}_{3d}\|_{1}+\|\hat{\mathbf{J}}_{2d}-\mathbf{J}_{2d}\|_{1}.$ (2)
Parameter-level supervision.
Since the proposed dataset 4DHuman-SKEL provides reliable pose $\hat{\boldsymbol{\theta}}$ and shape $\hat{\boldsymbol{\beta}}$ parameters, we impose supervision directly on these values:
$\mathcal{L}_{\beta}=\|\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}\|_{1},\quad\mathcal{L}_{\theta}=\|\hat{\boldsymbol{\theta}}-\boldsymbol{\theta}\|_{1}.$ (3)
Figure 3: Visual comparison between the proposed SKEL-CF and HSMR [39]. Our proposed SKEL-CF achieves more precise skeletal estimations (best viewed in zoom). Additional visual results are provided in the supplementary material.
Coarse-to-fine refinement.
SKEL-CF adopts a coarse-to-fine refinement strategy. The encoder first produces an initial estimate of the parameters, denoted as $\boldsymbol{\Theta}_{0}=\{\boldsymbol{\theta}_{0},\boldsymbol{\beta}_{0},\boldsymbol{\pi}_{0}\}$. The decoder then refines these predictions across $L$ layers, where the $i$-th layer outputs $\boldsymbol{\Theta}_{i}=\{\boldsymbol{\theta}_{i},\boldsymbol{\beta}_{i},\boldsymbol{\pi}_{i}\}$. The final layer produces the refined prediction $\boldsymbol{\Theta}_{L}$. Both the encoder output and the final decoder output are optimized using the same objective function $\mathcal{L}_{\text{skel}}$:
$\mathcal{L}_{\text{enc}}=\mathcal{L}_{\text{skel}},\quad\mathcal{L}_{\text{dec}}=\mathcal{L}_{\text{skel}}.$ (4)
Iterative refinement.
To stabilize training, we adopt an iterative refinement strategy with layer-wise supervision. Specifically, each decoder layer predicts a residual update to the parameters based on the output of the previous layer, progressively refining the estimate. To facilitate effective learning at each stage, we apply auxiliary supervision [44, 3] to every decoder layer using the same refinement loss.
$\mathcal{L}_{\text{refine}}=\sum_{i=1}^{L}\|\hat{\boldsymbol{\theta}}_{i}-\boldsymbol{\theta}_{i}\|_{1},$ (5)
where $L$ denotes the total number of decoder layers, and $\hat{\boldsymbol{\theta}}_{i}$ denotes the corresponding ground-truth pose parameters. We observe that optimizing $\mathcal{L}_{\text{skel}}$ provides similar performance. Since computing 2D keypoint predictions via the SKEL forward process is computationally expensive, we restrict optimization to the pose parameters.
Note that coarse-to-fine strategy, auxiliary supervision, and iterative refinement are widely adopted in the DETR-based object detection frameworks [44, 3]. The experiments to demonstrate the effectiveness of iterative refinement are provided in Appendix S2.
Overall optimization goal.
The overall loss function combines the above components for both the coarse encoder prediction and the iterative refinement in decoder:
$\mathcal{L}_{\text{tot}}=\mathcal{L}_{\text{dec}}+\mathcal{L}_{\text{enc}}+\lambda_{ref}\mathcal{L}_{\text{refine}},$ (6)
where $\lambda_{ref}$ is the coefficient for the refinement loss.
Implementation details.
Following HSMR [39], we adopt ViTPose-H [41], pretrained on COCO [22], as the encoder backbone. The hyper-parameters are set as $\lambda_{kp}=0.05$, $\lambda_{\beta}=0.0005$, $\lambda_{\theta}=0.001$, and $\lambda_{ref}=0.1$. More implementation details are provided in Appendix S3.
4 Experiment
In this section, we first introduce the datasets and evaluation metrics in Section 4.1. We then provide a comparison between SKEL-based methods in Section 4.2 and SMPL-based approaches in Section 4.3. Finally, we present the ablation studies and additional discussions in Section 4.4.
Method 3DPW [35] EMDB [14] SPEC-SYN [17] MOYO [33] MOYO-HARD
[…]
SMPL [23]-based Approaches
SPEC [17] 96.5 53.2 118.5 138.9 87.7 161.3 83.5 56.9 98.9 - - - - - -
CLIFF [21] 69.0 43.0 81.2 103.5 68.3 123.7 128.5 55.8 139.0 154.6 109.3 155.7 - - -
HMR2.0a [6] 69.8 44.4 82.2 97.8 61.5 120.0 133.3 55.8 153.0 - - - - - -
TokenHMR [4] 70.5 43.8 86.0 88.1 49.8 104.2 110.5 51.8 127.6 - - - - - -
WHAM [30] 57.8 35.9 68.7 79.7 50.4 94.4 - - - - - - - - -
ReFit [36] 57.6 38.2 67.6 91.7 55.5 106.2 103.6 51.3 116.3 - - - - - -
CLIFF [21] 72.0 46.6 85.0 97.1 61.3 113.2 109.9 55.6 124.6 - - - - - -
HMR2.0b [6] 81.3 54.3 93.1 118.5 79.2 140.6 150.7 67.6 172.9 123.3 90.4 142.2 - - -
CameraHMR [27] 62.7 38.7 73.4 73.2 43.9 85.6 66.0 37.0 79.1 85.1 50.2 95.3 96.0 61.9 111.8
SKEL [15]-based Approaches
HSMR [39] 81.5 54.8 - - - - - - - 104.5 79.6 120.1 120.0 97.7 140.5
SKEL-CF (Ours) 61.5 38.7 73.5 72.0 44.5 84.7 69.4 37.1 83.4 85.0 51.4 91.9 90.0 61.5 102.5
Table 2: Quantitative comparison of mesh recovery using the SMPL [23] human model on 3DPW [35], EMDB [14], SPEC-SYN [17], MOYO [33], and the more challenging MOYO-HARD dataset. The proposed SKEL-based method, SKEL-CF, achieves performance competitive with CameraHMR [27], with more pronounced improvements on complex motion scenarios as MOYO-HARD.
Figure 4: Visual comparison between the proposed SKEL-CF and CameraHMR [27]. The proposed SKEL-CF is built upon the SKEL [15] representation, which enforces anatomically consistent joint motion, resulting in more natural pose predictions compared to the SMPL [23]-based CameraHMR [27] (best viewed in zoom). Additional visual examples are provided in the supplementary material.
4.1 Datasets and evaluation metrics
Evaluated datasets.
We evaluate SKEL-CF on five representative datasets covering diverse environments and motion complexities, following standard practice [27, 39, 11, 4]. 3DPW [35] provides in-the-wild videos with accurate 3D annotations captured by moving cameras. Human3.6M [9] includes controlled indoor actions (e.g., sitting, walking, greeting) with motion-capture ground truth. EMDB [14] offers high-quality pose and shape sequences captured with electromagnetic sensors and handheld cameras. SPEC-SYN [17] is synthetic with explicit camera intrinsics and diverse perspectives for controlled camera evaluation. In contrast, MOYO [33] contains large-scale yoga sequences with extreme poses, frequent self-occlusions, and ground contact, providing the most challenging and diverse test scenario. Since most yoga videos in MOYO starts and ends with the same static poses, we select the front-view camera among the eight available viewpoints and remove the first and last 25% of frames to retain the complex and diverse motion segments, forming the curated MOYO-HARD subset. More details are provided in Appendix, Figure S1., where we additionally include visualizations of the removed first and last 25% segments of the MOYO videos.
Evaluation metrics.
We evaluate the 3D human body recovery using MPJPE, PA-MPJPE, and PVE. Among them, MPJPE and PA-MPJPE serve as sparse evaluations, measuring the Euclidean distance between the predicted and ground-truth kinematic joints before and after rigid alignment, respectively. In contrast, PVE provides a dense evaluation, computing the vertex-level error between the predicted and ground-truth meshes, thereby offering a more comprehensive assessment of surface reconstruction accuracy. The 2D keypoint metrics, PCK@0.05 and PCK@0.1, which quantify the ratio of projected keypoints within a normalized distance threshold to the ground truth.
4.2 Comparison with SKEL-based approaches.
Quantitative results.
We compare SKEL-based methods in Table 1. Among end-to-end approaches, SKEL-CF consistently outperforms prior work HSMR [39] across all datasets. On the challenging MOYO [33] benchmark, SKEL-CF achieves up to 18.6% MPJPE and 35.4% PA-MPJPE improvements, demonstrating strong robustness under large pose variations and occlusions. We also evaluate two-stage baseline, HMR2.0 + SKEL fit and CameraHMR + SKEL fit. While CameraHMR + SKEL fit achieves competitive results on MOYO and MOYO-HARD, its performance drops on 3DPW, indicating limited stability across datasets. Moreover, the fitting-based approach is computationally expensive (approximately 4 min/img on RTX 3090), whereas SKEL-CF runs in a single forward pass (approximately 0.8 s/img), making it both more efficient and more reliable.
Visual results.
We provide a visual comparison with HSMR [39] in Figure 3. As shown, the proposed SKEL-CF achieves more accurate skeletal estimation and mesh reconstruction. More visual examples similar to Figure 3 are included in Appendix, Figure S3.
Figure 5: Illustration of the mesh refinement. The light green meshes represent the initial coarse estimations, while the dark green meshes denote the progressively refined final results.
4.3 Comparison with SMPL-based approaches
Quantitative results.
We further compare SKEL-CF with recent state-of-the-art SMPL-based human mesh recovery methods across four representative datasets: 3DPW [35], EMDB [14], SPEC-SYN [17], MOYO [33] and more challenging MOYO-HARD, as summarized in Table 2. Note that SMPL [23] is less constrained compared to SKEL [15], enable achieving strong quantitative performance but occasionally producing anatomical inconsistency or unnatural joint motions(see Figure 1). Despite this inherent difference, SKEL-CF achieves comparable performance to CameraHMR [27], the strongest SMPL-based model trained on 4DHuman, demonstrating that kinematically constrained SKEL representations can reach a similar level of numerical precision while maintaining stronger physical plausibility.
Notably, on the challenging MOYO [33] dataset, which contains complex yoga sequences with strong self-occlusion and body–ground interactions, SKEL-CF achieves state-of-the-art performance in dense mesh prediction (PVE) and overall reconstruction accuracy. Furthermore, on the MOYO-HARD subset, focused on dynamic and highly articulated poses, SKEL-CF consistently outperforms all SMPL-based baselines across all metrics (MPJPE, PA-MPJPE, and PVE). This highlights its superior ability to maintain accurate and physically realistic motion representations under extreme articulations. These results demonstrate that SKEL-CF provides a scalable and anatomically faithful framework for human motion analysis. For completeness, we also report the 2D keypoint results in Table 3. Our approach achieves comparable accuracy under PCK@0.1, indicating that the projected mesh provides a consistent and accurate fit to the 2D image plane.
Visual results.
We provide a visual comparison with the SMPL [23]-based state-of-the-art method CameraHMR [27] in Figure 4. As shown, the proposed SKEL-CF produces anatomically more consistent joint motions, benefiting from the structural constraints of the SKEL [15] model. In contrast, CameraHMR tends to generate unnatural joint configurations under complex motion scenarios. Additional visualizations similar to Figure 4 are provided in Appendix, Figure S2.
4.4 Discussion
Ablation study.
To analyze the contribution of each component, we conduct an ablation study on MOYO [33] and COCO [22], representing 3D and 2D evaluation metrics, respectively. We start from a baseline that removes camera modeling, coarse-to-fine (C2F), and iterative refinement, which is architecturally comparable to HSMR [39]. In this controlled setting, the only difference from HSMR is the training data. The large performance gap on MOYO (e.g., PA-MPJPE: 53.7 vs. 79.6) highlights the effectiveness of our higher-quality training annotations. Building on this baseline, we progressively activate three key components: an explicit camera model, the coarse-to-fine strategy, and the iterative refinement stage, and report the corresponding performance under both 3D and 2D metrics. Incorporating the camera model notably improves 3D reconstruction accuracy. The coarse-to-fine strategy further enhances the 2D metrics. Finally, the iterative refinement stage brings consistent improvements across both 3D and 2D evaluations.
Iterative mesh refinement.
We further visualize the iterative mesh refinement process of SKEL-CF in Figure 5. This visualization clearly demonstrates the effectiveness of the proposed coarse-to-fine refinement strategy. Additional visual results illustrating the refinement process are provided in Appendix S2.
Method PCK@0.05 $\uparrow$ PCK@0.1 $\uparrow$
PyMAF [43] 0.68 0.86
CLIFF [21] 0.64 0.88
PARE [16] 0.72 0.91
PyMAF-X [42] 0.79 0.93
HMR2.0a [6] 0.79 0.94
CameraHMR [27] 0.84 0.94
HSMR [39] 0.85 0.96
SKEL-CF (Ours) 0.80 0.93
Table 3: 2D Keypoint Performance on COCO [22]. For completeness, we also report PCK scores of the projected keypoints at thresholds of 0.05 and 0.1, showing that SKEL-CF achieves comparable 2D keypoint quality while maintaining significant better biomechanical mesh quality.
Components MOYO [33] COCO [22]
[…]
88.8 53.7 97.0 0.76 0.91
89.5 54.8 97.9 0.79 0.93
84.7 50.6 91.7 0.67 0.91
86.2 52.1 93.5 0.77 0.92
85.0 51.4 91.9 0.80 0.93
Table 4: Ablation study on MOYO [33] and COCO [22]. Cam denotes the camera model, C2F represents the coarse-to-fine strategy, and Refine indicates the iterative refinement stage.
5 Conclusion
In this work, we introduced SKEL-CF, a biomechanical skeleton and surface mesh recovery pipeline built upon the SKEL representation. By transforming the high-quality SMPL dataset into a SKEL-based format, termed 4DHuman-SKEL, we established a strong foundation for accurate and generalizable training. Our integration of an explicit camera model effectively mitigates the limitations of weak-perspective assumptions seen in previous works such as HSMR, enabling robust pose estimation under diverse viewpoints. Furthermore, the proposed coarse-to-fine refinement strategy enables stable and high-fidelity reconstruction results. Overall, SKEL-CF sets a new baseline for biomechanical skeleton and surface recovery, providing a practical and extensible framework for future research in human modeling and motion understanding.
Appendix
S1 Organization
This appendix document is organized as follows:
  • Additional video demonstrations showcasing both our SKEL-CF visualizations and comparative results with HSMR [39] and CameraHMR [27] are provided in the project page: https://pokerman8.github.io/SKEL-CF/
  • Details on the iterative refinement are in Section S2, more visual results illustrating our coarse-to-fine refinement strategy are shown in Fig. S4, as a supplement to Fig. 5.
  • Implementation details are in Section S3.
  • Additional qualitative results comparing our method with the SKEL-based model HSMR [39] are presented in Fig. S3.
  • Additional qualitative comparisons with the SMPL [23]-based approach CameraHMR [27] are provided in Fig S2.
  • Details on the MOYO-HARD subset are in Section S4.
  • Section S7 analyzes the per-layer attention behavior of Pose, Beta, and Cam tokens.
S2 Details of iterative refinement
Figure S4 presents additional visual examples from LSP-ext [10], highlighting how the refinement process progressively improves pose and shape estimation. We further validate this design in Table S1: the proposed sparse refinement supervision consistently improves performance on MOYO and MOYO-HARD and achieves accuracy comparable to full per-layer prediction, while being substantially more efficient. In addition, repeating a single decoder layer degrades performance, especially on MOYO-HARD, highlighting the importance of progressive refinement across layers.
Method Loss Decoder Layer 3DPW MOYO MOYO-HARD
[…]
Sparse super. (Ours) $\mathcal{L}_{refine}$ 6 $\times$ 1 61.5 38.7 73.5 85.0 51.4 91.9 90.0 61.5 102.5
Full predict. $\mathcal{L}_{skel}$ 6 $\times$ 1 61.3 38.6 73.3 86.4 53.2 93.9 92.1 65.2 105.6
One layer $\mathcal{L}_{refine}$ 1 $\times$ 6 62.5 40.1 74.8 86.8 54.4 94.2 96.8 70.1 111.5
Table S1: Further study of iterative refinement. Sparse super. applies refinement supervision across decoder layers, Full predict. predicts full SKEL parameters at every layer, and One layer repeats a single layer for multiple refinement steps. “Decoder Layer $\times$ Refine Time” summarizes the refinement configuration.
Table S2: COCO results under different camera extrinsic factorizations ($s$, $t_{x}$, $t_{y}$). (A) Fix final SKEL [15] parameters and $s$, vary layer-wise $(t_{x},t_{y})$. (B) Fix final SKEL parameters and $(t_{x},t_{y})$, vary layer-wise $s$.
Layer (A) Final Pose + $s$; Layer $(t_{x},t_{y})$ (B) Final Pose + $(t_{x},t_{y})$; Layer $s$
PCK@0.05 PCK@0.1 PCK@0.05 PCK@0.1
0 0.1147 0.4156 0.0164 0.0597
1 0.0462 0.2493 0.0171 0.0604
2 0.0312 0.1876 0.0193 0.0669
3 0.1205 0.6088 0.0290 0.1038
4 0.5592 0.8837 0.0689 0.2643
5 0.8037 0.9361 0.8037 0.9360
S3 Implementation details
Following HSMR [39], we adopt ViTPose-H [41], pretrained on COCO [22], as the encoder backbone. ViTPose-H consists of 32 transformer layers, each with 16 attention heads and a hidden dimension of 1280. The decoder is a standard transformer decoder with $L=6$ layers, consistent with the design in HSMR [39]. We train SKEL-CF using the AdamW [24] optimizer with $\beta_{1}$, $\beta_{2}$ setting to 0.9, 0.999, weight decay setting to $1\times 10^{-4}$, a batch size of 64 and a learning rate of $1\times 10^{-5}$, preceded by a one-epoch warm-up. Training is conducted for 30 epochs on 8 NVIDIA A100 GPUs, taking approximately 120 hours in total. The hyper-parameters are set as $\lambda_{kp}=0.05$, $\lambda_{\beta}=0.0005$, $\lambda_{\theta}=0.001$, and $\lambda_{ref}=0.1$. Notably, our training schedule is significantly shorter than that of HSMR [39], which trains for 100 epochs.
S4 Details of MOYO-HARD dataset
The MOYO [33] test set consists of 29 yoga sequences, each recorded simultaneously by eight cameras capturing the same pose from different viewpoints. Each sequence progresses from simple poses to complex ones and then returns to simpler poses, as shown in Figure S1. To specifically analyze performance under challenging poses, we focus on the middle portion of each sequence, where the motions are most complex. Accordingly, we construct the MOYO-HARD subset by selecting the middle 50% of frames from videos captured by Camera #1 (front view), resulting in a total of 9,725 images.
Figure S1: Pose Sequence from MOYO. The video starts and ends with static, low-motion poses, while the middle segment contains diverse, high-complexity movements, reflecting the overall difficulty distribution of the dataset.
S5 Visual comparison with SMPLs
To more comprehensively evaluate our model, we present in Fig. S2 additional visual comparisons extending the results shown in Fig. 4 of the main paper. These examples further compare our method with the state-of-the-art SMPL [23]-based approach, CameraHMR [27]. It indicates that our method can reconstruct more natural pose with accurate anatomical joints.
S6 Visual comparison with SKELs
To further compare SKEL [15]-based approaches, Fig. S3 shows additional qualitative comparisons with the state-of-the-art HSMR [39] pipeline. Unlike HSMR, our method incorporates explicit camera modeling, enabling more reliable and accurate pose reconstruction under varying viewpoints.
S7 Per-Layer Attention Analysis
Figure S5 visualizes the attention regions of the Pose, Beta, and Cam tokens across different decoder layers. In Figure S5(a), we show representative examples of the final Pose, Beta, and Cam tokens. We observe that the Pose token primarily attends to the interior body regions, the Beta token focuses on the body silhouette, and the Cam token concentrates on the surrounding environment, as it must aggregate contextual information to estimate the camera extrinsics. In Figure S5(b)–(c), we visualize layer-wise attention for the same sample. The Pose token gradually shifts across different body parts, the Beta token evolves from the head region to the full silhouette, and the Cam token expands from the human region to the environment for estimating camera extrinsics. As shown in Table S2, our ablation further reveals that layers 1–5 mainly learn 2D translation, while the final layer predominantly captures scale (depth).
Figure S2: Additional comparisons with SMPL [23]-based models. Compared with the state-of-the-art SMPL-based method CameraHMR [27], our approach produces more natural and anatomically consistent joint predictions.
Figure S3: Additional comparisons with SKEL [15]-based models. Compared with the state-of-the-art HSMR [39] method, our approach predicts human poses that are more realistic and accurate.
Figure S4: Additional iterative refinement results on LSP-ext [10]. These examples illustrate how our iterative refinement strategy progressively improves pose and shape estimation, yielding more accurate and stable reconstructions across diverse human poses.
Figure S5: Layer-wise attention analysis of Pose, Beta, and Cam tokens. The Pose token progressively shifts across body parts, the Beta token expands from head to full silhouette, and the Cam token gradually incorporates scene context for estimating camera extrinsics.